Bexley:

The circumference of one coin is 8.03 cm. The circumference of another coin is 0.33 cm smaller. Find the diameter of the smaller coin and use 22/7 for pie

BabyBB2:

Based on the given information, we can solve for the diameter of the smaller coin. Let's start by finding the circumference of the smaller coin. We know that it is 0.33 cm smaller than the first coin, which has a circumference of 8.03 cm. So, the circumference of the smaller coin is: 8.03 cm - 0.33 cm = 7.70 cm Using the formula for circumference, which is C = pi x d, where C is the circumference and d is the diameter, we can solve for the diameter of the smaller coin. 7.70 cm = (22/7) x d d = 7.70 cm x (7/22) d = 2.45 cm Therefore, the diameter of the smaller coin is approximately 2.45 cm.